OpenStudy (anonymous):

(1 pt) What is the equation of the graphed line? A. B. C. D.

OpenStudy (igreen):

Just post the graph, I don't need the options.

OpenStudy (igreen):

Okay, I think I need the options..xD

OpenStudy (igreen):

Well it's going to be something like.. y = 0.16x

OpenStudy (anonymous):

LOLz but here they come

OpenStudy (anonymous):

A y=6x B y =1/6 x c y = -1/6 x d y = -6x

OpenStudy (igreen):

C. \(\dfrac{1}{6} \approx 0.16\)

OpenStudy (anonymous):

so negative

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

No, it's not a negative. The answer is C.

OpenStudy (igreen):

Sorry B.

OpenStudy (igreen):

If it was a negative, the line would be flipped over.